OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 B. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 D. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3

OpenStudy (anonymous):

someone please explain this to me

OpenStudy (anonymous):

Ok so slope-intercept form is represented by the equation y = mx + b In this equation, the 'm' is your slope, and the 'b' is your y-intercept.

OpenStudy (anonymous):

So let's get your equation: -11x + 9y = -12 into y = mx + b form.

OpenStudy (anonymous):

im like so confused on it i do not understand i am not getting the right answers when i work it out

OpenStudy (anonymous):